OpenStudy (anonymous):

a rectangular shape with an area of 192 square feet. The dimensions are x^2 feet wide by x^2– 61 feet long. How wide is the shape? a. 3 feet c. 64 feet b. 8 feet d. 131 feet

OpenStudy (anonymous):

\[192=(x^2)*(x^2-61)\]\[192=x^4-61x^2\]\[x^4-61x^2-192=0\]\[(x-8)(x+8)(x^2+3)=0\] x=8 \[8^2=64ft\]\[8^2-61=3ft\]If you like my answer ,click good answer,TY

OpenStudy (anonymous):

a. 3 feet

OpenStudy (anonymous):

a. 3 feet